# Multi-Agent Collaborative Framework for Intelligent IT Operations: An AOI System with Context-Aware Compression and Dynamic Task Scheduling

### v1

@@ III-A Problem statement and framework overview
To tackle these challenges, we designed an innovative multi-agent framework integrated with large language model (LLM) reasoning and hierarchical memory management. Our solution consists of the following components: Multi-agent collaborative architecture- This module comprises three specialized agents—Observer for task decomposition and coordination, Probe for safe read-only information gathering, and Executor for risk-controlled system modification—ensuring clear responsibility separation and operational safety. LLM-based context compressor- This module employs a sliding-window mechanism and domain-aware summarization to intelligently compress operational context, achieving high compression ratios while retaining critical diagnostic and remediation information. Dynamic task scheduling strategy- This module adaptively prioritizes probing and execution tasks based on real-time system states, balancing exploration (information collection) and exploitation (fault remediation) to optimize mean time to resolution (MTTR) and task success rate. Three-layer memory architecture- This module includes Raw Context Storage for unprocessed operational data, Task Queue Management for structured execution instructions, and Compressed Context Cache for LLM-processed strategic information, enabling efficient context management across multi-scale operations.
III-B Problem Formulation
Let $\mathcal{S}=\{s_{1},s_{2},...,s_{n}\}$ represent the set of system components under management, where each component $s_{i}$ has a state vector $\mathbf{x}_{i}(t)\in\mathbb{R}^{d}$ at time $t$. Given a user task $\tau$ and current system state $\mathbf{X}(t)=[\mathbf{x}_{1}(t),\mathbf{x}_{2}(t),...,\mathbf{x}_{n}(t)]^{T}$, our objective is to find an optimal sequence of actions $\mathcal{A}=\{a_{1},a_{2},...,a_{k}\}$ that transforms the system to a desired target state $\mathbf{X}^{*}$ while minimizing the operational cost function:
$\mathcal{J}(\mathcal{A})=\alpha\cdot T_{completion}+\beta\cdot C_{resource}+\gamma\cdot R_{risk}$ (1)
where $T_{completion}$ is the Mean Time to Resolution (MTTR), $C_{resource}$ represents computational and network overhead, $R_{risk}$ quantifies the operational risk score of potential system instability or SLA violations, and $\alpha$, $\beta$, $\gamma$ are weighting parameters.
III-C AOI Framework Architecture
Figure 2 illustrates the overall architecture of our AOI framework, consisting of four core components and a three-layer memory management system.
III-C1 Observer Agent
The Observer Agent serves as the central coordination unit responsible for task analysis, decomposition, and strategic decision-making. Its core functionalities include:
Task Decomposition Algorithm: Given a complex task $\tau$, the Observer employs a hierarchical decomposition strategy:
@@ III-C1 Observer Agent
1:Task $\tau$, system state $\mathbf{X}(t)$, context $C_{\text{compressed}}$
2:Subtask queue $\mathcal{Q}_{\text{subtasks}}$
3:Initialize $\mathcal{Q}_{\text{subtasks}}\leftarrow\emptyset$
4:Analyze task complexity $\mathcal{C}(\tau)$
5:if $\mathcal{C}(\tau)>\theta_{\text{complex}}$ then
6:  Decompose $\tau$ into atomic subtasks $\{\tau_{1},\tau_{2},\dots,\tau_{m}\}$
7:  for all $\tau_{i}$ in decomposed tasks do
8:   Determine task type: $\text{type}(\tau_{i})\in\{\text{probe},\text{execute}\}$
9:   Estimate resource requirements $R(\tau_{i})$
10:   Assign priority $P(\tau_{i})$ based on dependency analysis
11:   Enqueue $(\tau_{i},\text{type}(\tau_{i}),R(\tau_{i}),P(\tau_{i}))$ to $\mathcal{Q}_{\text{subtasks}}$
12:  end for
13:else
14:  Enqueue $\tau$ directly to $\mathcal{Q}_{\text{subtasks}}$
15:end ifreturn $\mathcal{Q}_{\text{subtasks}}$
@@
$Schedule(t)=\arg\max_{a\in\{probe,execute\}}\mathbb{E}[Reward(a,\mathbf{X}(t),C(t))]$ (2)
where the expected reward function considers information gain for probing actions and progress toward goal state for execution actions. Specifically, the reward for a probe action is quantified by the expected reduction in system state uncertainty (information gain), while the reward for an execute action is based on the estimated progress towards the target system state $\mathbf{X}^{*}$.
III-C2 Probe Agent
The Probe Agent is designed as a strictly read-only component that safely gathers system information without modifying system state. Key design principles include:
Safety-by-Design Principle: The Probe Agent operates under the core principle of least privilege, executing strictly in a read-only mode to eliminate any risk of unintended system modification during diagnostic phases.
Safety Constraints: The Probe Agent operates under strict safety rules: - Whitelist of allowed operations: {SELECT, SHOW, DESCRIBE, GET, LIST, READ} - Blacklist of forbidden operations: {DELETE, UPDATE, INSERT, DROP, ALTER, TRUNCATE} - File operations limited to read-only modes - Network operations restricted to query-type requests
Information Collection Strategy: The Probe Agent employs a structured approach to information gathering:
@@ III-C2 Probe Agent
1:Probe task $\tau_{\text{probe}}$, target system $\mathcal{S}$
2:Raw information $I_{\text{raw}}$
3:Initialize $I_{\text{raw}}\leftarrow\emptyset$
4:Generate probe script $script\leftarrow\textsc{GenerateScript}(\tau_{\text{probe}})$
5:Validate script safety: $\textsc{ValidateSafety}(script)$
6:if IsSafe(script) then
7:  for all command cmdcmd in scriptscript do
8:   $result\leftarrow\textsc{ExecuteReadOnly}(cmd,\mathcal{S})$
9:   Append resultresult to $I_{\text{raw}}$
10:   if ErrorOccurred(result) then
11:     Log error and continue with next command
12:   end if
13:  end for
14:else
15:  Reject unsafe script and request revision
16:end if
17:Store $I_{\text{raw}}$ in memory–raw context return $I_{\text{raw}}$
III-C3 Executor Agent
The Executor Agent is responsible for implementing system modifications to achieve desired operational outcomes. Unlike the Probe Agent, it can modify system state but operates under careful coordination with the Observer.
Conservative Execution Strategy: The Executor follows a "cautiously optimistic" execution approach. To ensure operational safety, every execution plan is preceded by a full-system checkpoint, enabling instantaneous rollback should any action yield an unexpected or critical failure, thereby implementing a fail-fast and recover-quickly paradigm.
Execution Strategy: The Executor follows a conservative execution approach:
@@ III-C3 Executor Agent
1:Execution task $\tau_{\text{exec}}$, target system $\mathcal{S}$
2:Execution result $R_{\text{exec}}$
3:Create checkpoint $checkpoint\leftarrow\textsc{CreateCheckpoint}(\mathcal{S})$
4:Generate execution plan $plan\leftarrow\textsc{GeneratePlan}(\tau_{\text{exec}})$ $\triangleright$ Plan generation includes pre-execution validation and dependency checks
5:for all action $a$ in planplan do
6:  if RequiresSystemInfo(a) then
7:   Query probe agent for current state
8:   Update action parameters based on current state
9:  end if
10:  $result\leftarrow\textsc{ExecuteAction}(a,\mathcal{S})$
11:  Validate execution result
12:  if CriticalFailureDetected(result) then
13:   Initiate rollback to checkpointcheckpoint
14:   break
15:  end if
16:end for
17:Store execution results in memory–raw context return $R_{\text{exec}}$
III-C4 Context Compressor
The Context Compressor leverages large language models to intelligently process and compress operational context while preserving critical information.
Semantic-Aware Compression: Unlike traditional keyword-based or statistical compression methods, our LLM-based compressor leverages the model’s semantic understanding capabilities to identify and preserve operationally critical information (e.g., error patterns, resource thresholds, causal relationships) while discarding redundant or low-signal data.
Sliding Window Compression: We implement a sliding window mechanism with 50% overlap to ensure context continuity and prevent the loss of critical information that might span the boundary between two consecutive windows:
$Window_{i}=[start_{i},start_{i}+window_{size}]$ (3)
where $start_{i}=i\times(window_{size}\times overlap_{ratio})$ and $overlap_{ratio}=0.5$.
Intelligent Compression Strategy: The compression algorithm prioritizes information based on criticality:
@@ III-C4 Context Compressor
1:Raw context $C_{\text{raw}}$, compression target ratio $r_{\text{target}}$
2:Compressed context $C_{\text{merged}}$
3:Split $C_{\text{raw}}$ into sliding windows $\{W_{1},W_{2},\dots,W_{k}\}$
4:Initialize $C_{\text{compressed}}\leftarrow\emptyset$
5:for all window $W_{i}$ do
6:  $I_{\text{critical}}\leftarrow\mathrm{ExtractCritical}(W_{i})$ $\triangleright$ LLM extracts fault signatures, error codes, and performance anomalies
7:  $S_{i}\leftarrow\mathrm{LLMSummarize}(W_{i},I_{\text{critical}})$
8:  Append $S_{i}$ to $C_{\text{compressed}}$
9:end for
[…]
11:if $\mathrm{Ratio}(C_{\text{merged}},C_{\text{raw}})>r_{\text{target}}$ then
12:  $C_{\text{merged}}\leftarrow\mathrm{SecondaryCompress}(C_{\text{merged}},r_{\text{target}})$
13:end ifreturn $C_{\text{merged}}$
III-D Three-Layer Memory Architecture
@@
Layer 1: Raw Context Storage - Stores unprocessed results from Probe and Executor agents with 24-hour retention. (Feeds into Context Compressor)
Layer 2: Task Queue Management - Maintains structured execution instructions with lifecycle management. (Managed by Observer and executed by Executor)
Layer 3: Compressed Context Cache - Houses LLM-processed information with 7-day retention for strategic decision-making. (Populated by Context Compressor and queried by Observer)
III-E Complexity Analysis
@@
IV Experiments
IV-A Experimental Setup
@@
All experiments were conducted in a controlled cloud-based environment designed to emulate realistic large-scale IT operations, with the infrastructure carefully configured to ensure reproducibility and enable controlled fault injection for stress testing. The experimental cluster consisted of eight AWS EC2 c5.4xlarge instances, each equipped with 16 vCPUs and 32 GB of RAM, supported by 1 TB of high-IOPS AWS EBS storage for persistent data management. All nodes were interconnected through a 10 Gbps network with adjustable latency injection to simulate diverse communication conditions. The software stack was built on Ubuntu 20.04 LTS, running Docker 20.10 and Kubernetes 1.24 for container orchestration. System monitoring and telemetry were implemented via an integrated observability framework composed of Prometheus, Grafana, and the ELK suite. The large language model (LLM) backend employed a fine-tuned GPT-4 API, specifically optimized for operations-related context understanding and summarization, ensuring accurate interpretation and efficient processing of system logs and alerts.
IV-B Datasets and Scenarios
@@
We evaluated the proposed AOI framework using two complementary datasets encompassing both synthetic and real-world operational environments. The first, the AIOpsLab Simulation Environment, is a synthetic dataset specifically designed for controlled experimentation, consisting of 1,000 operational scenarios that span 50 distinct fault types. It covers a wide range of conditions, from isolated single-service failures to complex cascading faults, each annotated by domain experts with corresponding ground-truth resolution procedures and expected post-recovery system states. The second, the Real-world Log Corpus (Loghub Benchmark), was employed to assess the framework’s practical applicability. This dataset aggregates production logs from diverse domains, including distributed storage, microservices, and enterprise-scale systems. To ensure compatibility with the AOI task structure, the logs were preprocessed through anonymization, message parsing, and scenario segmentation. Across both datasets, the evaluation encompassed four representative categories of IT operational events: (1) Service Failure Recovery, including database connectivity disruptions, web server crashes, and microservice communication breakdowns; (2) Performance Degradation, characterized by CPU and memory bottlenecks, network latency spikes, and storage I/O constraints; (3) Configuration Drift, involving version mismatches, dependency conflicts, and runtime misconfigurations; and (4) Security Incidents, covering unauthorized access attempts, certificate expirations, and policy enforcement violations.
IV-C Baseline Methods
@@
To comprehensively evaluate the effectiveness of the proposed AOI framework, we compared it against four representative and reproducible baseline approaches that capture the major paradigms in automated IT operations. The Rule-based Expert System (RES) represents traditional automation techniques relying on deterministic rules implemented through Ansible playbooks and custom operational scripts. The Traditional AIOps Pipeline (TAP) follows a classical machine learning architecture that integrates anomaly detection modules with predefined scripted responses for fault remediation. The Single-Agent LLM (SA-LLM) baseline applies a large language model directly to operational decision-making tasks without incorporating multi-agent collaboration or contextual memory mechanisms. Finally, the Baseline Multi-Agent System (B-MAS) extends conventional automation through agent-based coordination but omits the intelligent context compression and adaptive scheduling strategies introduced in our AOI framework.
IV-D Evaluation Metrics
@@
IV-E Implementation Details
The system is architected for reproducibility and robustness, with each agent implemented as a containerized microservice using Python 3.9 and the FastAPI framework to handle inter-agent communication. For context compression, we integrate the OpenAI GPT-4 API, employing custom prompts optimized for the operational context; this process exhibits an average API latency of 2.3 seconds per compression operation. Memory management is handled by a Redis Cluster, which provides distributed storage, automatic failover, and data persistence. To ensure operational safety, critical system properties are formally verified using TLA+ specifications, complemented by runtime monitoring that enforces a set of custom safety rules.
V Results
V-A Main Experimental Results
As presented in Table II, our AOI framework demonstrates superior performance across all evaluation metrics compared to the baseline methods. It achieves the highest task success rate of 94.2%, marking a 34.9% and 9.0% improvement over the traditional rule-based and the best multi-agent baselines, respectively. Furthermore, AOI reduces the mean time to resolution (MTTR) to 22.1 minutes, a 34.4% reduction from the best baseline of 33.7 minutes. The framework also attains a context compression ratio (CCR) of 72.4% while maintaining a high information preservation score (IPS) of 92.8%, validating the effectiveness of our LLM-based compression. Finally, AOI upholds robust safety and accuracy, evidenced by the lowest false positive rate (3.1%) and the highest system safety score (96.7%).
TABLE II: Main Experimental Results Comparison

### v2

@@ 4.1 Problem statement and framework overview
To tackle these challenges, we designed an innovative multi-agent framework integrated with large language model (LLM) reasoning and hierarchical memory management. Our solution consists of the following components: Multi-agent collaborative architecture- This module comprises three specialized agents—Observer for task decomposition and coordination, Probe for safe read-only information gathering, and Executor for risk-controlled system modification—ensuring clear responsibility separation and operational safety. LLM-based context compressor- This module employs a sliding-window mechanism and domain-aware summarization to intelligently compress operational context, achieving high compression ratios while retaining critical diagnostic and remediation information. Dynamic task scheduling strategy- This module adaptively prioritizes probing and execution tasks based on real-time system states, balancing exploration (information collection) and exploitation (fault remediation) to optimize mean time to resolution (MTTR) and task success rate. Three-layer memory architecture- This module comprises Raw Context Storage for unprocessed data, Task Queue Management for structured execution, and a Compressed Context Cache for LLM-processed strategic information, enabling efficient multi-scale context management.
4.2 Problem Formulation
[…]
4.3 AOI Framework Architecture
Figure 1 shows the AOI architecture, comprising 4 core components and a 3-layer memory system.
4.3.1 Observer Agent
The Observer Agent serves as the central coordination unit responsible for task analysis, decomposition, and strategic decision-making. Its core functionalities include: Task Decomposition Algorithm: Given a complex task $\tau$, the Observer employs a hierarchical decomposition strategy shown in Algorithm 1.
@@ 4.3.1 Observer Agent
0: Task $\tau$, System State $\mathbf{X}(t)$, Context $C_{compressed}$
0: Subtask Queue $\mathcal{Q}_{subtasks}$
1: Initialize $\mathcal{Q}_{subtasks}\leftarrow\emptyset$
2: Analyze task complexity $\mathcal{C}(\tau)$
3:if $\mathcal{C}(\tau)>\theta_{complex}$ then
4:  Decompose $\tau$ into atomic subtasks $\{\tau_{1},\tau_{2},...,\tau_{m}\}$
5:  for each $\tau_{i}$ in decomposed tasks do
6:   Determine task type: $type(\tau_{i})\in\{probe,execute\}$
7:   Estimate resource requirements $R(\tau_{i})$
8:   Assign priority $P(\tau_{i})$ based on dependency analysis
9:   Enqueue $(\tau_{i},type(\tau_{i}),R(\tau_{i}),P(\tau_{i}))$ to $\mathcal{Q}_{subtasks}$
10:  end for
11:else
12:  Enqueue $\tau$ directly to $\mathcal{Q}_{subtasks}$
13:end if
14:return $\mathcal{Q}_{subtasks}$
@@
$S(t)=\arg\max_{a\in\{\mathrm{probe},\mathrm{execute}\}}\mathbb{E}[\mathrm{Reward}(a,\mathbf{X}(t),C(t))]$
where S means Schedule, the expected reward function considers information gain for probing actions and progress toward goal state for execution actions. Specifically, the reward for a probe action is quantified by the expected reduction in system state uncertainty (information gain), while the reward for an execute action is based on the estimated progress towards the target system state $\mathbf{X}^{*}$.
4.3.2 Probe Agent
The Probe Agent is a strictly read-only component for safe system information gathering. Safety-by-Design: It follows the least-privilege principle to prevent unintended modifications during diagnostics. Safety Constraints: The agent is limited to non-destructive query and read operations (e.g., database introspection and log/file reads), while any state-changing actions (e.g., writes, deletions, schema changes, or destructive commands) are explicitly prohibited. File access remains read-only, and network interactions are restricted to query-type requests.
Information Collection Strategy: The Probe employs a structured probing strategy (Algorithm 2).
@@ 4.3.2 Probe Agent
0: Probe Task $\tau_{probe}$, Target System $\mathcal{S}$
0: Raw Information $I_{raw}$
1: Initialize $I_{raw}\leftarrow\emptyset$
2: Generate probe script $script\leftarrow GenerateScript(\tau_{probe})$
3: Validate script safety: $ValidateSafety(script)$
4:if script is safe then
5:  for each command cmdcmd in script do
6:   Execute $result\leftarrow ExecuteReadOnly(cmd,\mathcal{S})$
7:   Append resultresult to $I_{raw}$
8:   if error occurred then
9:    Log error and continue with next command
10:   end if
11:  end for
12:else
13:  Reject unsafe script and request revision
14:end if
15: Store $I_{raw}$ in Memory-Raw Context
16:return $I_{raw}$
4.3.3 Executor Agent
The Executor Agent implements system modifications to achieve target operational outcomes. Unlike the Probe Agent, it can modify system state under coordination with the Observer. Conservative Execution Strategy: The Executor adopts a cautiously optimistic approach. Each execution is preceded by a full-system checkpoint, enabling instant rollback upon unexpected or critical failures, thereby supporting a fail-fast, recover-quickly paradigm (Algorithm 3).
@@ 4.3.3 Executor Agent
0: Execution Task $\tau_{exec}$, Target System $\mathcal{S}$
0: Execution Result $R_{exec}$
1: Create checkpoint $checkpoint\leftarrow CreateCheckpoint(\mathcal{S})$
2: Generate execution plan $plan\leftarrow GeneratePlan(\tau_{exec})$ {Plan generation includes pre-execution validation and dependency checks}
3:for each action $a$ in plan do
4:  if RequiresSystemInfo(a) then
5:   Query Probe Agent for current state
6:   Update action parameters based on current state
7:  end if
8:  Execute $result\leftarrow ExecuteAction(a,\mathcal{S})$
9:  Validate execution result
10:  if critical failure detected then
11:   Initiate rollback to checkpointcheckpoint
12:   BREAK
13:  end if
14:end for
15: Store execution results in Memory-Raw Context
16:return $R_{exec}$
4.3.4 Context Compressor
The Context Compressor leverages large language models to intelligently process and compress operational context while preserving critical information. Semantic-Aware Compression: Unlike traditional keyword-based or statistical compression methods, our LLM-based compressor leverages the model’s semantic understanding capabilities to identify and preserve operationally critical information (e.g., error patterns, resource thresholds, causal relationships) while discarding redundant or low-signal data. Sliding Window Compression: We implement a sliding window mechanism with 50% overlap to ensure context continuity and prevent the loss of critical information that might span the boundary between two consecutive windows: $\mathrm{Window_{i}}=[start_{i},start_{i}+window_{size}]$ where $start_{i}=i\times(window_{size}\times overlap_{ratio})$ and $overlap_{ratio}=0.5$. Intelligent Compression Strategy: The compression algorithm prioritizes information based on criticality:
@@ 4.3.4 Context Compressor
0: Raw Context $C_{raw}$, Compression Target Ratio $r_{target}$
0: Compressed Context $C_{compressed}$
1: Split $C_{raw}$ into sliding windows $\{W_{1},W_{2},...,W_{k}\}$
2: Initialize $C_{compressed}\leftarrow\emptyset$
3:for each window $W_{i}$ do
4:  Identify critical information: $I_{critical}\leftarrow\mathrm{ExtractCritical}(W_{i})$ {Using LLM to extract fault signatures, error codes, and performance anomalies}
5:  Generate summary: $S_{i}\leftarrow\mathrm{LLMSummarize}(W_{i},I_{critical})$
6:  Append $S_{i}$ to $C_{compressed}$
7:end for
8: Merge overlapping summaries: $C_{merged}\leftarrow\mathrm{MergeOverlaps}(C_{compressed})$
9: Apply secondary compression if needed to meet $r_{target}$
10:return $C_{merged}$
4.4 Three-Layer Memory Architecture
@@
Layer 1: Raw Context Storage — Stores unprocessed results from Probe and Executor agents with 24-hour retention. (Feeds into Context Compressor). Layer 2: Task Queue Management — Maintains structured execution instructions with lifecycle management. (Managed by Observer and executed by Executor). Layer 3: Compressed Context Cache — Houses LLM-processed information with 7-day retention for strategic decision-making. (Populated by Context Compressor and queried by Observer).
4.5 Complexity Analysis
@@
5 Experiments
5.1 Experimental Setup
@@
All experiments were conducted in a controlled cloud-based environment designed to emulate realistic large-scale IT operations, with the infrastructure carefully configured to ensure reproducibility and enable controlled fault injection for stress testing. The experimental cluster consisted of eight AWS EC2 c5.4xlarge instances, each equipped with 16 vCPUs and 32 GB of RAM, supported by high-IOPS persistent storage. All nodes were interconnected through a 10 Gbps network with adjustable latency injection to simulate diverse communication conditions. The software stack was built on Ubuntu 20.04 LTS, running Docker and Kubernetes for container orchestration. System monitoring and telemetry were implemented via an integrated observability framework composed of Prometheus, Grafana, and the ELK stack Prometheus Authors (2024); Grafana Labs (2024); Elastic (2024). The large language model (LLM) backend employed a fine-tuned GPT-4 API optimized for operations-related context understanding and summarization, enabling accurate interpretation and efficient processing of system logs and alerts OpenAI (2023).
5.2 Datasets and Scenarios
@@
We evaluate AOI on two complementary datasets spanning synthetic and real-world environments. The AIOpsLab Simulation Environment contains 1,000 fault-injection scenarios covering 50 fault types, ranging from isolated failures to cascading faults, each annotated with ground-truth resolutions and post-recovery states Chen et al. (2025). To assess practical applicability, we additionally use the Loghub Benchmark, which aggregates real-world production logs from distributed storage systems, microservices, and enterprise infrastructures Zhu et al. (2023). Logs are preprocessed via anonymization, parsing, and scenario segmentation to align with AOI’s task structure. Across both datasets, evaluation focuses on four representative operational event categories: service failure recovery, performance degradation, configuration drift, and security incidents.
5.3 Baseline Methods
@@
To comprehensively evaluate the effectiveness of the proposed AOI framework, we compared it against four representative and reproducible baseline approaches that capture the major paradigms in automated IT operations. The Rule-based Expert System (RES) Xu et al. (2020); Sarazin et al. (2021) represents traditional automation techniques relying on deterministic rules implemented through Ansible playbooks and custom operational scripts. The Traditional AIOps Pipeline (TAP) Dang et al. (2019) follows a classical machine learning architecture that integrates anomaly detection modules with predefined scripted responses for fault remediation. The Single-Agent LLM (SA-LLM) baseline Jiang et al. (2023) applies a large language model directly to operational decision-making tasks without incorporating multi-agent collaboration or contextual memory mechanisms. Finally, the Baseline Multi-Agent System (B-MAS) D’Aniello et al. (2021) extends conventional automation through agent-based coordination but omits the intelligent context compression and adaptive scheduling strategies introduced in our AOI framework.
5.4 Evaluation Metrics
@@
5.5 Implementation Details
The system is designed for reproducibility and robustness. Each agent is implemented as a containerized microservice in Python 3.9 using FastAPI for inter-agent communication FastAPI Contributors (2024). Context compression integrates the OpenAI GPT-4 API with custom prompts optimized for operational contexts, incurring an average latency of 2.3 seconds per compression OpenAI (2023). Memory management relies on a Redis Cluster for distributed storage, failover, and persistence. Operational safety is ensured via formal TLA+ specifications and runtime monitoring with custom safety rules Lamport, Leslie (2023).
6 Results
6.1 Main Experimental Results
As shown in Table 2, AOI outperforms all baselines across metrics. It achieves the highest task success rate (94.2%), improving by 34.9% and 9.0% over rule-based and best multi-agent baselines, respectively. AOI reduces MTTR to 22.1 minutes, a 34.4% reduction from the best baseline (33.7 minutes). It also attains a 72.4% context compression ratio with 92.8% information preservation, validating the LLM-based compression. Finally, AOI achieves the lowest false positive rate (3.1%) and the highest safety score (96.7%).
Table 2: Main Experimental Results Comparison